OpenStudy (anonymous):

Which of the following is a polynomial with roots -2, -3i, and 3i ?

OpenStudy (campbell_st):

well with roots of 3i, -3i and -2you can have a polynomial P(x) = (x -3i)(x + 3i)(x + 2) just distribute and the first 2 binomial factors will be the difference of 2 squares

OpenStudy (anonymous):

-3i?

OpenStudy (campbell_st):

its a complex number solution.... what you need to know is \[i^2 = -1\] so \[3i \times -3i = -9i^2\] now subsitute \[i^2 = -1 \] and its \[-9 \times -1 = 9\]

OpenStudy (anonymous):

x3 + 4x2 + 9x + 24 x3 - 4x2 + 9x - 24 x3 + 2x2 + 9x + 18 x3 - 2x2 + 9x - 18

OpenStudy (campbell_st):

well not quite if you start with the 1st 2 binomials P(x) = (x - 3i)(x + 3i)(x + 2) (x - 3i)(x + 3i) = x^2 -3xi + 3xi -9i^2 which I've simplified below \[P(x) = (x^2 + 9)(x + 2)\]

OpenStudy (campbell_st):

you just need to distribute to get your polynomial

OpenStudy (anonymous):

oh ok